OpenStudy (anonymous):

ƒ(x ) = 3x + 1 and ƒ -1 = x-1 (they are both over) /3

OpenStudy (anonymous):

@ilovebmth1234

OpenStudy (anonymous):

these are your options 22 2 -2

RhondaSommer (rhondasommer):

can you screen shot it?

OpenStudy (anonymous):

i'll try drawing it or sum lol

RhondaSommer (rhondasommer):

thanks. the waty you wrote it out didnt make much sense

OpenStudy (ilovebmth1234):

im not sure xC

OpenStudy (anonymous):

\[ƒ(x ) = 3x + 1 and ƒ -1 = \frac{ x-1 }{ 3 } then ƒ -1(7) =\]

OpenStudy (anonymous):

i thought i could do the words ;-; and it is fine @ilovebmth1234

OpenStudy (anonymous):

lol its kinda hard idk how to screenshot on laptops

RhondaSommer (rhondasommer):

is is \[f-1x-1\] or is it \[(f-1)/3=(x-1)/3\]

RhondaSommer (rhondasommer):

what kind of laptop is it?

OpenStudy (anonymous):

thinkpad for alternative

OpenStudy (anonymous):

did that work?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i'm a just guess

RhondaSommer (rhondasommer):

Use The Keyboard Shortcut - Alt + PrtScn. In Windows 8.1 and Windows 10, you can also take screenshots of the active window. Open the window you want to capture and press Alt + PrtScn on your keyboard